OpenStudy (anonymous):

Find f(-7) if f(x) = 2x + 8. -6 2 3 22

OpenStudy (jdoe0001):

ok, start at the beginning and finish when you're done :)

OpenStudy (anonymous):

f(-7) = -14 + 8 f(-7) = -6

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

\[\huge\color{magenta}{f(x) =2x+8 }\] \[\huge\color{magenta}{f(-7) =2(-7)+8 = -14+8 =-6 } \]

OpenStudy (anonymous):

^^^YOU ARE SO FANCY. :3

OpenStudy (anonymous):

yeaa :D

OpenStudy (anonymous):

i am flattered